OpenStudy (anonymous):

Which of the following constants can be added to x^2 - x to form a perfect square trinomial? 1/4 1/2 1

OpenStudy (lgbasallote):

what is the coefficient of -x?

OpenStudy (anonymous):

1

OpenStudy (lgbasallote):

it's -1...divide it by 2 what do you gET/

OpenStudy (anonymous):

but -1 isn't a perfect square

OpenStudy (anonymous):

0.5

OpenStudy (lgbasallote):

well it's not yet done lo

OpenStudy (lgbasallote):

square 0.5

OpenStudy (anonymous):

it's 1/4 1/2=.5 .5^2=.25 or 1/4

OpenStudy (anonymous):

what one

OpenStudy (lgbasallote):

0.25 = 1/4 1/4 is the answer

OpenStudy (anonymous):

ok thank you